OpenStudy (anonymous):

An airplane is flying at an elevation of 1,326 meters when it begins a steady descent. If it descends 61 meters per minute, how long will it take the airplane to reach an elevation of 350 meters?

hero (hero):

You only need to post it one time

OpenStudy (anonymous):

Srry!

hero (hero):

Find the distance between 1,326 and 350, then whatever that is, divide it by 61

hero (hero):

so, 350 + x = 1,326 x = 1326 - 350 x = 976 976/61 = ?

OpenStudy (anonymous):

16 thx

hero (hero):

16 minutes

OpenStudy (anonymous):

thx

hero (hero):

So the plane descended 976 meters and it took 16 minutes to do so.